# 3D Feature Distillation with Object-Centric Priors

### v3

@@ 3.3 View-Independent Feature Distillation
Even though the above feature-cloud $Z^{3D}$ could be directly used for open-vocabulary grounding in 3D, its construction is computationally intensive and requires a lot of expensive resources, such as access to multiple camera views, view-aligned 2D instance segmentation masks, as well as a set of text descriptions to compute informativeness metrics. Such utilities are rarely available in open-ended scenarios, especially in robotic applications, where usually only single-view RGB-D images from sensors mounted on the robot are provided. To tackle this, we wish to distill all the above knowledge from the feature-cloud $Z^{3D}$ into a 3D encoder that receives only a partial point-cloud from single-view posed RGB-D. Hence, the only assumption that we make during inference is access to camera intrinsic and extrinsic parameters, which is a mild requirement in most robotic works.
In particular, given a partial colored point-cloud from view $v$: $P_{v}\in\mathbb{R}^{M_{v}\times 6}$ (3D coordinates plus colors), we train a 3D encoder $\mathcal{E}_{\theta}:\mathbb{R}^{M_{v}\times 6}\rightarrow\mathbb{R}^{M_{v}\times C}$ such that $\mathcal{E}_{\theta}(P_{v})=Z^{3D}$. Notice that the distillation target $Z^{3D}$ is independent of view $v$. Following  [12, 15] we use cosine distance loss:
@@
See Appendix B.2 for training implementation details. With such a setup, we can obtain 3D features that: (i) are co-embedded in CLIP text space, so they can be leveraged for 3D segmentation tasks from open-vocabulary queries via computing cosine similarities between CLIP text embeddings $Q$ and the predicted feature cloud: $\hat{S}_{i}=\texttt{argmax}_{i}\;\texttt{cos}(\hat{\mathbf{z}}_{i}^{3D},Q)$, (ii) are ensured to be optimally informative per object, due to the usage of the semantic informativeness metric to compute $Z^{3D}$, (iii) maintain 3D spatial consistency in object boundaries, due to performing object-wise instead of point-wise fusion when computing $Z^{3D}$, and (iv) are encouraged to be view-independent, as the same features $Z^{3D}$ are utilized as distillation targets regardless of the input view $v$. Importantly, no labels, prompts, or segmentation masks are needed at test-time to reproduce the fused feature-cloud, while obtaining it amounts to a single forward pass of our 3D encoder, hence offering real-time performance.
Figure 4: Open-Vocabulary 3D Referring Segmentation. Examples of learned 3D features and grounding heatmaps from open-ended language queries (class names, attributes, user affordances, and open instance-specific concepts) in scenes from MV-TOD dataset. Points are colored based on their query similarity (higher towards red). We note that table points are excluded from similarity computation in our visualizations.
@@ 4 Experiments
In our experiments, we explore the following questions: (i) Sec. 4.1: What are the contributions of our proposed object-centric priors for multi-view feature fusion? Does the dense number of views of our proposed dataset also contribute? (ii) Sec. 4.2: How does our method compare to previous open-vocabulary approaches for 3D semantic and referring segmentation tasks? Are the learned features robust to open-ended language? (iii) Sec. 4.3: What are the generalization capabilities of our learned 3D representation in novel domains and novel tasks (3D instance segmentation)? (iv) Sec. 4.4: Can we leverage our 3D learned representation for language-guided 6-DoF robotic grasping?
@@ 4.1 Multi-view Feature Fusion Ablation Studies
To evaluate the contributions of our proposed object-centric priors, we conduct ablation studies on the multi-view feature fusion pipeline, where we compare 3D referring segmentation results of obtained 3D features in held-out scenes of MV-TOD. We highlight that here we aim to establish a performance upper bound that the feature fusion method can provide for distillation, and not the distilled features themselves.
Figure 5: Referring segmentation accuracy (Pr@25 (%)) vs. number of utilized views.
We ablate: (i) patch-wise vs. object-wise fusion, (ii) MaskCLIP [8] patch-level vs. CLIP [7] masked crop features, (iii) inclusion of visibility ($\Lambda_{v,i}$) and semantic informativeness ($G_{v,i}$) metrics for view selection. Results in Table 2.
Effect of object-centric priors We observe that all components contribute positively to the quality of the 3D features. Our proposed $G_{v,i}$ metric boosts mIoU across both point- and object-wise fusion ($57.0\%$ vs. $44.2\%$ and $83.1\%$ vs. $65.6\%$ respectively). Further, we observe that the usage of spatial priors for object-wise fusion and object-level features leads to both higher segmentation crispness ($25.7\%$ mIoU delta), as well as higher grounding precision ($42.5\%$ Pr@75 delta). See qualitative comparisons in Appendix D.
Effect of the number of views We ablate the 3D referring segmentation performance based on the number of input views in Fig. 5, where novel viewpoints are added incrementally. We observe that in both setups (point- and object-wise) fusing features from more views leads to improvements, with a small plateauing behavior around 40 views. We believe this is an encouraging result for leveraging dense multi-view coverage in feature distillation pipelines, as we propose with the introduction of MV-TOD.
4.2 Open-Vocabulary 3D Segmentation Results
@@
mIoU Pr@25 Pr@50 Pr@75 mIoU mAcc
OpenScene  [12] 73 29.32 44.00 24.51 11.26 21.79 32.14
OpenMask3D  [6] 73 65.38 73.05 63.99 57.40 59.47 66.48
DROP-CLIP (Ours) 73 82.67 86.11 82.43 79.23 75.41 80.02
DROP-CLIP (Ours) 73 66.56 75.73 67.55 59.88 62.04 70.74
OpenSeg→3D  [9] 1 12.89 17.36 2.38 0.23 12.83 17.21
MaskCLIP→3D  [8] 1 25.64 40.36 18.69 6.95 20.97 32.09
DROP-CLIP (Ours) 1 62.31 71.96 62.75 53.85 54.48 64.41
@@
Table 3: Referring and Semantic segmentation results on MV-TOD test split. Methods with denote upper-bound 3D features, whereas DROP-CLIP denotes our distilled model. Methods with →3D produce 2D predictions that are projected to 3D to compute metrics. Method with * denotes further usage of ground-truth segmentation masks.
In this section, we compare referring and semantic segmentation performance of our distilled features vs. previous open-vocabulary approaches, both in multi-view and in single-view settings. For multi-view, we compare our trained model with OpenScene [12] and OpenMask3D [6] methods, where the full point-cloud from all $73$ views is given as input.
Figure 6: Referring segmentation accuracy (Pr@25 (%)) vs. different language query types.
@@
We note that for these baselines we obtain the upper-bound 3D features as before, as we observed that our trained model already outperforms them, so we refrained from also distilling features from baselines (details in Appendix C2). For single-view, we feed our network with partial point-cloud from projected RGB-D pair, and compare with 2D baselines MaskCLIP [8] and OpenSeg [9]. Our model slightly outperforms the OpenMask3D upper bound baseline in the multi-view setting ($+1.18\%$ in referring and $+2.57\%$ in semantic segmentation), while significantly outperforming 2D baselines in the single-view setting ($>30\%$ in both tasks). Importantly, single-view results closely match the multi-view ones ($\sim-4.0\%$), suggesting that DROP-CLIP indeed learns view-independent features.
@@
Open-ended queries We evaluate the robustness of our model in different types of input language queries, organized in 4 families (class name - e.g. “cereal box", class + attribute - e.g. “brown cereal box", open - e.g. “chocolate Kellogs", and affordance - e.g. “I want something sweet‘). Comparative results are presented in Fig. 6 and qualitative in Fig. 4. We observe that our method achieves high grounding accuracy in all query types, even when using single-view.
4.3 Zero-Shot Transfer to Novel Domains / Tasks
Method OCID-VLG  [31] REGRAD  [30]
IoU Pr@25 IoU Pr@25
MaskCLIP→3D  [8] 24.1 30.9 33.2 39.0
DROP-CLIP (Ours) 46.2 48.9 59.1 63.0
Table 4: Referring segmentation results in OCID-VLG [31] and REGRAD [30] datasets
Generalization to Novel Domains We evaluate the 3D referring segmentation performance of our trained model when applied zero-shot in novel tabletop domains. We test in 500 scenes from OCID-VLG [31] using the dataset’s instance-wise open queries, as well as in 1000 scenes from REGRAD [30], using each model’s class name as a query. Only single-view input is provided for both datasets.
Method $mIoU$ $AP_{25}$ $AP_{50}$
SAM  [39] 70.11 95.26 79.88
DROP-CLIP (S) 80.83 91.92 86.83
Mask3D  [40] 14.41 18.65 3.41
DROP-CLIP (F) 88.37 93.13 91.47
@@ 4.3 Zero-Shot Transfer to Novel Domains / Tasks
Table 5: Zero-shot 3D instance segmentation results in MV-TOD.
We compare with MaskCLIP [8] as above and report results in Table 4. We note that test datasets contain both novel object instances (REGRAD) and classes (OCID-VLG). We observe that our method provides a significant performance boost across both domains ($22.1\%$ mIoU delta in OCID-VLG and $25.9\%$ in REGRAD).
Zero-Shot 3D Instance Segmentation Since our method has been distilled from features with object-level priors, we demonstrate that it can be used out-of-the-box for 3D instance segmentation, via clustering the 3D features (see Appendix E for implementation details). We report results in MV-TOD in Table 5, where we compare with SAM [39] with single-view images, as well as Mask3D  [41] with full point-clouds (transferred from ScanRefer [1] with room layout). Mask3D struggles to generalize to tabletop domains, whereas our method achieves comparable performance with SAM for segmenting from single-view, even without being explicitly trained for instance segmentation.
4.4 Open-Vocabulary Language-guided Robotic Grasping
@@
Figure 7: Language-guided 6-DoF grasping trial with real robot (left), 3D features, grounding and grasp proposals (right).
@@
In this section, we wish to illustrate the applicability of DROP-CLIP in a language-guided robotic grasping scenario. We integrate our method with a 6-DoF grasp detection network [42], to segment and then propose gripper poses for picking a target object indicated verbally. We randomly place 5-12 objects on a tabletop with different levels of clutter, and query the robot to pick the target object and place it in a fixed position. The user instruction is open-vocabulary and can involve open object descriptions, attributes, or affordances. We conducted 50 trials in Gazebo [43] and 10 with a real robot, and observed grounding accuracy of 84% and 80% respectively, and a final success rate of 64% and 60%, where failures were mostly due to grasp proposals that are outside of the robot’s kinematic range or motion planning that lead to a collision with other objects and the table. Our setup and example trials are shown in Fig. 7, while more details and qualitative results are provided in Appendix E. A video of robot demonstrations is provided as supplementary material.

### v4

@@ 3.3 View-Independent Feature Distillation
Even though the above feature-cloud $Z^{3D}$ could be directly used for open-vocabulary grounding in 3D, its construction is computationally intensive and requires a lot of expensive resources, such as access to multiple camera views, view-aligned 2D instance segmentation masks, as well as textual prompts to compute informativeness metrics. Such utilities are rarely available in open-ended scenarios, especially in robotic applications, where usually only single-view RGB-D images from sensors mounted on the robot are provided. To tackle this, we wish to distill all the above knowledge from the feature-cloud $Z^{3D}$ with an encoder network that receives only a partial point-cloud from single-view posed RGB-D. Hence, the only assumption that we make during inference is access to camera intrinsic and extrinsic parameters, which is a mild requirement in most robotic pipelines.
In particular, given a partial colored point-cloud from view $v$: $P_{v}\in\mathbb{R}^{M_{v}\times 6}$, we train an encoder $\mathcal{E}_{\theta}:\mathbb{R}^{M_{v}\times 6}\rightarrow\mathbb{R}^{M_{v}\times C}$ such that $\mathcal{E}_{\theta}(P_{v})=Z^{3D}$. Notice that the distillation target $Z^{3D}$ is independent of view $v$. Following  [12, 15] we use cosine distance loss:
@@
See Appendix A.2 for training implementation details. With such a setup, we can obtain 3D features that: (i) are co-embedded in CLIP text space, so they can be leveraged for 3D segmentation tasks from open-vocabulary queries, (ii) are ensured to be optimally informative per object, due to the usage of the semantic informativeness metric to compute $Z^{3D}$, (iii) maintain 3D spatial consistency in object boundaries, due to performing object-wise instead of point-wise fusion when computing $Z^{3D}$, and (iv) are encouraged to be view-independent, as the same features $Z^{3D}$ are utilized as distillation targets regardless of the input view $v$. Importantly, no labels, prompts, or segmentation masks are needed at test-time to reproduce the fused feature-cloud, while obtaining it amounts to a single forward pass of our 3D encoder, hence offering real-time performance.
3.4 Open-Vocabulary 3D Segmentation
Given a predicted feature-cloud $\hat{Z}^{3D}=\mathcal{E}_{\theta}(P_{v})$, we can perform 3D grounding tasks from open-vocabularies by computing cosine similarities between CLIP text embeddings and $\hat{Z}^{3D}$.
Semantic segmentation In this task, the queries correspond to an open-set of textual prompts $Q=\{\mathbf{q}_{k}\}_{k=1}^{\mathcal{K}}$ describing $\mathcal{K}$ semantic classes. A class for each point $\hat{Y}\in\{1,\dots,\mathcal{K}\}^{M}$ is given by : $\hat{Y}=\texttt{argmax}_{k}\;\texttt{cos}(\hat{Z}^{3D},\mathbf{q}_{k})$.
Referring segmentation Here the user provides an open-vocabulary query $\mathbf{q}^{+}$ referring to a particular object instance, and optionally a set of negative prompts $Q^{-}\in\mathbb{R}^{N^{-}\times C}$, which in practise can be initialized from an open-set as above or with canonical phrases (e.g. ‘object’, ‘thing’ etc.) [13]. Similarity scores are converted to probabilities: $\mathcal{P}=\texttt{softmax}\left(\frac{1}{\gamma}\cdot\texttt{cos}(\hat{Z}^{3D},\,[\textbf{q}^{+},Q^{-}]^{T})\right)$, where $\gamma$ a temperature hyper-parameter and $\mathcal{P}=[\boldsymbol{\rho}^{+},\mathcal{P}^{-}]$ probabilities of positive matching $\boldsymbol{\rho}^{+}\in\mathbb{R}^{M}$ and negative matching $\mathcal{P}^{-}=[\boldsymbol{\rho}^{-}_{1},\dots,\boldsymbol{\rho}^{-}_{N^{-}}]\in\mathbb{R}^{M\times N^{-}}$ respectively. The final 3D segmentation is given by $\hat{S}_{i}=\left(\boldsymbol{\rho}^{+}_{i}>\texttt{max}_{j}\,\mathcal{P}^{-}_{i,j}\right)$ , or by thresholding $\boldsymbol{\rho}^{+}$ with a fixed threshold $s_{thr}$ (see ablations in Appendix A.3)
Instance segmentation Since our encoder has been distilled with the aid of instance-wise segmentation masks, the obtained features can be utilized out-of-the-box for 3D instance segmentation tasks. We demonstrate that with a simple clustering algorithm over $\hat{Z}^{3D}$ we can obtain 3D instance segmentation masks for cluttered scenes, where naive 3D coordinate clustering would fail, performing competitively with popular segmentation methods in unseen data in the single-view setting (see Sec. 4.3). We refer the reader to Appendix A.6.2 for implementation details and related visualizations.
@@ 4 Experiments
Figure 4: Open-Vocabulary 3D Referring Segmentation in MV-TOD. Examples of learned 3D features and grounding heatmaps from open-ended language queries (class names, attributes, user affordances, and open instance-specific concepts) in scenes from MV-TOD dataset. Points are colored based on their query similarity (higher towards red). We note that table points are excluded from similarity computation in our visualizations.
We design our experiments to explore the following questions: (i) Sec. 4.1: What are the contributions of our proposed object-centric priors for multi-view feature fusion? Does the dense number of views of our proposed dataset also contribute? (ii) Sec. 4.2: How does our method compare to state-of-the-art open-vocabulary approaches for semantic and referring segmentation tasks, both in multi- and in single-view settings? Is it robust to open-ended language? (iii) Sec. 4.3: What are the zero-shot generalization capabilities of our learned 3D representation in novel datasets that contain real-world scenes, as well as for the novel task of 3D instance segmentation? (v) Sec. 4.4: Can we leverage DROP-CLIP for language-guided 6-DoF robotic grasping?
@@ 4.1 Multi-view Feature Fusion Ablation Studies
To evaluate the contributions of our proposed object-centric priors, we conduct ablation studies on the multi-view feature fusion pipeline, where we compare 3D referring segmentation results of obtained 3D features in held-out scenes of MV-TOD. We highlight that here we aim to establish a performance upper-bound that the feature fusion method can provide for distillation, and not the distilled features themselves. We ablate: (i) patch-wise vs. object-wise fusion, (ii) MaskCLIP [8] patch-level vs. CLIP [7] masked crop features, (iii) inclusion of visibility ($\Lambda_{v,i}$) and semantic informativeness ($G_{v,i}$) metrics for view selection. We report 3D segmentation metrics mIoU and Pr@X  [39]. Results in Table 2.
Figure 5: Referring segmentation precision vs. number of utilized views.
Effect of object-centric priors We observe that all components contribute positively to the quality of the 3D features. Our proposed $G_{v,i}$ metric boosts mIoU across both point- and object-wise fusion ($57.0\%$ vs. $44.2\%$ and $83.1\%$ vs. $65.6\%$ respectively). Further, we observe that the usage of spatial priors for object-wise fusion and object-level features leads to drastic improvements, both in segmentation crispness ($25.7\%$ mIoU delta), as well as in grounding precision ($42.5\%$ Pr@75 delta).
Effect of the number of views We ablate the 3D referring segmentation performance based on the number of input views in Fig. 5, where novel viewpoints are added incrementally. We observe that in both setups (point- and object-wise) fusing features from more views leads to improvements, with a small plateauing behavior around 40 views. We believe this is an encouraging result for leveraging dense multi-view coverage in feature distillation pipelines, as we propose with MV-TOD. Please see Appendix A.3 for extended ablation studies that justify the design choices behind our fusion strategy, and Appendix A.5 for qualitative comparisons with vanilla approaches.
4.2 Open-Vocabulary 3D Segmentation Results in MV-TOD
In this section, we compare referring and semantic segmentation performance of our distilled features vs. previous open-vocabulary approaches, both in multi-view and in single-view settings.
@@ 4.2 Open-Vocabulary 3D Segmentation Results in MV-TOD
mIoU Pr@25 Pr@50 Pr@75 mIoU mAcc25
OpenScene 73 29.3 44.0 24.5 11.3 21.8 32.1
OpenMask3D 73 65.4 73.1 64.0 57.4 59.5 66.5
DROP-CLIP 73 82.7 86.1 82.4 79.2 75.4 80.0
DROP-CLIP 73 66.6 75.7 67.6 59.9 62.0 70.7
OpenSeg→3D 1 12.9 17.4 2.4 0.2 12.8 17.2
MaskCLIP→3D 1 25.6 40.4 18.7 7.0 21.0 32.1
DROP-CLIP 1 62.3 72.0 62.8 53.9 54.5 64.4
@@
Table 3: Referring and Semantic segmentation results on MV-TOD test split. Methods with denote upper-bound 3D features, whereas DROP-CLIP denotes our distilled model. Methods with →3D produce 2D predictions that are projected to 3D to compute metrics. Methods with * denote further usage of ground-truth segmentation masks.
@@
For multi-view, we compare our trained model with OpenScene [12] and OpenMask3D [6] methods, where the full point-cloud from all $73$ views is given as input. We note that for these baselines we obtain the upper-bound 3D features as before, as we observed that our trained model already outperforms them, so we refrained from also distilling features from baselines. For single-view, we feed our network with partial point-cloud from projected RGB-D pair, and compare with 2D baselines MaskCLIP [8] and OpenSeg [9] (see implementation details in Appendix A.4). Our model slightly outperforms the OpenMask3D upper bound baseline in the multi-view setting ($+1.18\%$ in referring and $+2.57\%$ in semantic segmentation), while significantly outperforming 2D baselines in the single-view setting ($>30\%$ in both tasks). Importantly, single-view results closely match the multi-view ones ($\sim-4.0\%$), suggesting that DROP-CLIP indeed learns view-independent features. See Appendix A.5 for more qualitative comparisons with baselines.
Figure 6: Referring segmentation precision vs. language query types.
@@
Open-ended queries We evaluate the robustness of our model in different types of input language queries, organized in 4 families (class name - e.g. “cereal", class + attribute - e.g. “brown cereal box", open - e.g. “chocolate Kellogs", and affordance - e.g. “I want something sweet‘). Comparative results are presented in Fig. 6 and qualitative in Fig. 4. We observe that single-view performance closely follows that of upper-bound across query types, with multi-word affordance queries being the highest family of failures, potentially due to the "bag-of-words" behavior of CLIP text embeddings [16].
Figure 7: Zero-Shot 3D Semantic Segmentation in Real Scenes: Comparison of different referring segmentation models for five example cluttered indoor scenes from the OCID dataset. PCA features are displayed at pixel-level for 2D methods LSeg and MaskCLIP and in 3D for our point-cloud-based DROP-CLIP. Heatmaps from 2D models LSeg and MaskCLIP are projected to 3D for direct comparison with DROP-CLIP.
4.3 Generalization to Novel Domains / Tasks
Zero-shot transfer to real-world scenes In this section, we evaluate the zero-shot generalization capability of DROP-CLIP in real-world scenes that contain objects and vocabulary outside the MV-TOD distribution. We test in the validation split of the OCID-VLG [31] dataset, which contains 1249 queries from 165 unique cluttered tabletop scenes. We compare with state-of-the-art 2D CLIP-based baselines LSeg [10], OpenSeg [9] and MaskCLIP [8] for the semantic segmentation task in the single-view setting as before.
Method OCID-VLG
mIoU mAcc50 mAcc75
LSeg→3D 44.1 37.9 23.5
OpenSeg→3D 47.1 33.1 19.1
MaskCLIP→3D 57.1 59.4 31.0
DROP-CLIP 60.2 60.1 38.7
Table 4: Zero-shot semantic segmentation results (%) in the validation split of the OCID-VLG real-world dataset.
Results are presented in Table 4. We find that even though fine-tuned in real data, baselines LSeg and OpenSeg under-perform compared to both MaskCLIP and our DROP-CLIP with a margin of $>10\%$ mIoU, which we attribute to the distribution gap between the fine-tuning dataset ADE20K [40] and OCID scenes. These baselines tend to ground multiple regions in the scene, while MaskCLIP and DROP-CLIP provides tighter segmentations (see Fig. 7). When considering the stricter mAcc75 metric, our approach scores a delta of $7.7\%$ compared to MaskCLIP, suggesting a significant gain in grounding accuracy, especially in cases where the object is heavily occluded. Failures cases were observed in grounding objects that significantly vary in geometry and semantics from the MV-TOD catalog. Please see Appendix A.6.1 for further zero-shot referring segmentation experiments and more qualitative results.
Method OCID-VLG MV-TOD
mIoU AP25 mIoU AP25
SAM 60.1 95.3 70.1 95.2
DROP-CLIP (S) 50.9 68.0 80.8 91.9
Mask3D - - 14.4 18.7
DROP-CLIP (F) - - 88.3 93.3
@@ 4.3 Generalization to Novel Domains / Tasks
Table 5: Zero-shot 3D instance segmentation results in OCID-VLG (real-world) and our MV-TOD dataset.
Zero-shot 3D instance segmentation We evaluate the potential of DROP-CLIP for out-of-the-box 3D instance segmentation via clustering the predicted features (see details in Appendix A.6.2). We conduct experiments for both the multi-view setting in MV-TOD, where we compare with Mask3D [41] transferred from the ScanRefer [1] checkpoint provided by the authors, where we feed full point-clouds from 73 views, as well as in OCID-VLG, where we compare with SAM [42] ViT-L model with single-view images. Results are summarized in Table 5. We observe that Mask3D struggles to generalize to tabletop domains, as it has been trained in room layout data with mostly furniture object categories. DROP-CLIP achieves an AP25 of 93.3%, illustrating that the learned 3D features can provide near-perfect instance segmentation in-distribution, even without explicit fine-tuning. When moving out-of-distribution in the single-view setting, we observe that DROP-CLIP achieves mIoU that is competitive with foundation segmentation method SAM ($50.9\%$ vs. $60.1$%). Failure cases include heavily cluttered regions of similar objects with same texture (e.g. food boxes), for which DROP-CLIP assigns very similar features that are identified as a single cluster.
4.4 Application: Language-guided Robotic Grasping
@@
Figure 8: Language-guided 6-DoF grasping: Example robot trial (left), 3D features, grounding and grasp proposal (right).
@@
In this section, we wish to illustrate the applicability of DROP-CLIP in a language-guided robotic grasping scenario. We integrate our method with a 6-DoF grasp detection network [43], which proposes gripper poses for picking a target object segmented by DROP-CLIP. We randomly place 5-12 objects on a tabletop with different levels of clutter, and query the robot to pick a specific object, potentially amongst distractor objects of the same category. The user instruction is open-vocabulary and can involve open object descriptions, attributes, or user-affordances. We conducted 50 trials in Gazebo [44] and 10 with a real robot, and observed grounding accuracy of 84% and 80% respectively, and a final success rate of 64% and 60%. Motion failures were mostly due to grasp proposals for which the motion planning led to collisions. Similar to OCID, grounding failures were due to unseen query concepts and / or instances. Setup and example trials are shown in Fig. 8, more details and visualizations in Appendix A.7 and a video of robot demonstrations is provided as supplementary material.